OpenStudy (anonymous):

A country's population in 1994 was 140 million. In 1998 it was 146 million. Estimate the popluation in 2011 using the exponential growth formula. Round your answer to the nearest million?

OpenStudy (wolf1728):

Najia2000 - are you coming back?

OpenStudy (wolf1728):

rate = 10^(log[end amt/bgn amt]/time) -1 rate = 10^(log[146,000,000/140,000,000]/4) -1 rate = 10^(0.004556205) -1 rate = 1.0105462738 -1 rate = .0105462738 Ending Amount (2011) = Beginning Amount * (1+ rate)^time Ending Amount = 140,000,000 * (1.0105462738)^17 Ending Amount = 140,000,000 * 1.1952410089 Ending Amount = 167,333,741 http://www.1728.org/expgrwth.htm

OpenStudy (anonymous):

oh wow, Thank you!!!

OpenStudy (wolf1728):

Najia2000 - I'm glad you came back!